OpenStudy (sierra12):

The inflation rate of the U.S. dollar is 3.7 percent. What this means is that every year, prices increase by 3.7 percent. If a quart of oil cost $1.69 five years ago, what does it cost now?

OpenStudy (sierra12):

help now!

OpenStudy (texaschic101):

1.69 -- start 1.69 + .037(1.69) = 1.69 + 0.06253 = 1.752 rounds to 1.75 -- yr 1 1.75 + .037(1.75) = 1.75 + 0.06475 = 1.814 rounds to 1.81 -- yr 2 1.81 + .037(1.81) = 1.81 + 0.06697 = 1.876 rounds to 1.88 -- yr 3 1.88 + .037(1.88) = 1.88 + 0.06956 = 1.949 rounds to 1.95 -- yr 4 1.95 + .037(1.95) = 1.95 + 0.07215 = 2.022 rounds to 2.02 --- yr 5